OpenStudy (anonymous):

Write an absolute value equation that can be solved to find the two points on a number line that are 10 units away from 3. (Use x as the variable)

OpenStudy (anonymous):

What two numbers are 10 units away from 3? Okay, so let's go 10 units above 3, that's 13. How do we get another one? We have to go the other way, so 3-10 is -7 The two numbers are -7 and 13, so hopefully it'll help you understand the next part. So we have to now write an equation, with absolute values, where it will equal 3. | X - 3 | Let's test it out. |13 -3| = 10 |-7 - 3| = 10 Both will end up being positive, since it's in absolute values.

OpenStudy (anonymous):

Slight typo, we're writing an equal where it will equal 10* not 3*, since the distance is 10.

OpenStudy (anonymous):

so its |13-3|=10 and |-7-3|=10?

OpenStudy (anonymous):

Yes, once you plug it in, but it's just asking for the equation using an X, so the answer would just be | x - 3 | = 10

OpenStudy (anonymous):

and that what i do yay thanks (: